# Unitary normalizers in finite-dimensional inclusions

### v1

@@ 1 Introduction
Normalizers occupy a central position in the structure theory of operator algebras, reflecting internal symmetries of inclusions and governing their associated crossed-product and groupoid constructions. Their importance is already evident in Dixmier’s seminal classification of maximal abelian subalgebras (MASAs) into singular, semi-regular, and regular (Cartan) types, where regularity is characterized by generation of the ambient algebra by the normalizer. This perspective was further developed by Feldman and Moore [10], who showed that Cartan subalgebras give rise to measured equivalence relations, and later, Renault ([16]) discovered a $C^{*}$-algebraic analogue of this result in the language of groupoids.
In subfactor theory, normalizers play an important role (see, for instance,[12, 6, 2, 3, 15]). Regular subfactors form a distinguished class encompassing crossed-product inclusions by finite groups and are closely related to depth-two phenomena and quantum symmetry. In particular, finite-index regular subfactors are known to admit unitary Pimsner–Popa bases and to arise from actions of weak Hopf algebras or quantum groupoids (see [3, 9]). Related classification results have also been obtained in [4] for finite-index regular inclusions of simple unital $C^{*}$-algebras, where regularity forces depth two and a crossed-product-type structure.
@@
Our approach is matrix-theoretic in nature. Recall that any inclusion $\mathcal{B}\subset\mathcal{A}$ of finite-dimensional von Neumann algebras is completely determined, up to isomorphism, by its inclusion matrix together with the dimension vectors of the simple summands. We introduce a new combinatorial invariant of an inclusion, called the normalizer matrix, which captures the structure of the normalizer purely at the level of the inclusion matrix. This invariant isolates the precise combinatorial obstructions to regularity and allows us to characterize regular inclusions in terms of elementary matrix data.
@@
As a consequence, any regular inclusion is isomorphic to tensor product and direct sum of inclusions $M_{n}(\mathbb{C})\subset M_{n}(\mathbb{C}),$ $\mathbb{C}\subset\oplus_{i=0}^{s-1}M_{n}(\mathbb{C})$ and $\mathbb{C}^{n}\subset M_{n}(\mathbb{C})$.
A second motivation for this work comes from the theory of unitary orthonormal (Pimsner–Popa) bases. Such bases play a central role in subfactor theory and quantum information theory, where they are closely related to unitary error bases and to protocols such as teleportation and dense coding. While the existence of unitary orthonormal bases is automatic for several basic inclusions, a complete characterization for general finite-dimensional inclusions has remained open. In a seminal work, Werner [18] established a one-to-one correspondence between tight teleportation schemes for tripartite systems $M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ and orthonormal bases of unitaries in $M_{n}(\mathbb{C})$. This correspondence was subsequently extended by Conlon et al. [8] to inclusions of the form $N\subset M_{n}(\mathbb{C})$, demonstrating how unitary orthonormal bases on general multi-matrix algebras can be used to study quantum teleportation schemes and quantum graphs. In these settings, tight teleportation schemes are shown to arise from orthonormal unitary Pimsner–Popa bases, generalizing Werner’s original framework. In all such applications, it is essential that the unitary orthonormal basis be contained in the normalizer.
@@
Theorem B: (See Theorem 4.2, Corollary 4.3 and 4.4) Let $(\mathcal{B}\subset\mathcal{A},E)$ be an inclusion of finite-dimensional von Neumann algebras, where $E$ denotes the unique trace-preserving conditional expectation. Then the following statements are equivalent:
@@
Finally, we show that every regular inclusion has depth two, linking our results to (weak) Hopf-algebraic structures. More precisely, if the inclusion $\mathcal{B}\subset\mathcal{A}$ is regular, then the corresponding inclusion matrix has depth $2$ in the sense of [5]. Thus, our results may be viewed as a finite-dimensional analogue of the structure theory of regular inclusions developed in the subfactor and $C^{*}$-algebraic settings.
@@ 2 Preliminaries
Notation: The block diagonal form $\begin{bmatrix}A_{0}&&\\ &\ddots\\ &&&A_{n}\end{bmatrix}$ will be denoted by $\text{bl-diag}(A_{1},\cdots,A_{n})$.
@@
Consider the inclusions $\mathcal{B}_{1}\subseteq\mathcal{A}_{1}$ and $\mathcal{B}_{2}\subseteq\mathcal{A}_{2}$ with inclusion matrix $A_{1}$ and $A_{2}$ respectively. Then the inclusion matrix of the direct sum $\mathcal{B}_{1}\oplus\mathcal{B}_{2}\subseteq\mathcal{A}_{1}\oplus\mathcal{A}_{2}$ is given by
$\begin{bmatrix}A_{1}&0\\ 0&A_{2}\par\end{bmatrix}$
and the inclusion matrix of the tensor product $\mathcal{B}_{1}\otimes\mathcal{B}_{2}\subseteq\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is given by $A_{1}\otimes A_{2}$.
@@
[1] Let $(\mathcal{B}\subseteq\mathcal{A},E)$ be a subalgebra system, with
$Am^{\prime}=n^{\prime},$
where $A$ is the inclusion matrix, and $m^{\prime},n^{\prime}$ are dimension vectors. Suppose $(\mathcal{B}\subseteq\mathcal{A},E)$ admits a unitary orthonormal basis with $d$-unitary elements. Then
$A^{t}n^{\prime}=dm^{\prime},$
Consequently,
$A^{t}Am^{\prime}=dm^{\prime};~~AA^{t}n^{\prime}=dn^{\prime}.$
In particular, $m^{\prime},n^{\prime}$ are Perron-Frobenius eigenvectors of $A^{t}A$ and $AA^{t}$ respectively with eigenvalue $d$.
@@
where $\mathcal{U(A)}$ denotes the group of unitaries of $\mathcal{A}$. The inclusion $\mathcal{B}\subseteq\mathcal{A}$ is called regular if $\mathcal{A}=\text{Span}\;\mathcal{N}_{\mathcal{A}}(\mathcal{B})$.
Equivalently, the normalizer can also be defined as
@@
where $Ad_{U}(x)=UxU^{*}.$
@@
    Scalar Inclusion: $(\mathbb{C}\subset\oplus_{i=0}^{s-1}M_{n}(\mathbb{C}),E_{1})$ where $E_{1}$ is the unique conditional expectation preserving the state
    $\varphi(\oplus X_{i})=\frac{1}{\sum_{i=0}^{s-1}n_{i}^{2}}\sum_{i=0}^{s-1}n_{i}~\mbox{trace}~(X_{i}),~~X_{i}\in M_{n_{i}},1\leq i\leq(s-1).$ (2.5)
@@
Suppose $(\mathcal{B}_{i}\subseteq\mathcal{A}_{i},E_{i})$ for $i=1,2$ are two finite dimensional inclusions with $U$-property. Suppose $\{U_{1}(i),\ldots,U_{d_{i}}(i)\}$ is a unitary o.n.b. in
$\mathcal{N}_{\mathcal{A}_{i}}(\mathcal{B}_{i})$ for $(\mathcal{B}_{i},\mathcal{A}_{i},E_{i})$ for $i=1,2.$ Take $\mathcal{B}=\mathcal{B}_{1}\otimes\mathcal{B}_{2}$. Consider it as a subalgebra of $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ in the natural way with the inclusion matrix being the tensor product of inclusion matrices. Then it is seen easily that $E=E_{1}\otimes E_{2}$ is a conditional expectation map from $\mathcal{A}$ to $\mathcal{B}.$ For this inclusion, $\mathcal{B}_{1}\otimes\mathcal{B}_{2}\subseteq\mathcal{A}_{1}\otimes\mathcal{A}_{2},E_{1}\otimes E_{2})$, we can observe that $\{U_{j}(1)\otimes U_{k}(2):0\leq j\leq d_{1};0\leq k\leq d_{2}\}$ is a unitary orthonormal basis in $\mathcal{N}_{\mathcal{A}_{1}\otimes\mathcal{A}_{2}}(\mathcal{B}_{1}\otimes\mathcal{B}_{2})$ as $\mathcal{N}_{\mathcal{A}_{1}}(\mathcal{B}_{1})\otimes\mathcal{N}_{\mathcal{A}_{2}}(\mathcal{B}_{2})\subseteq\mathcal{N}_{\mathcal{A}_{1}\otimes\mathcal{A}_{2}}(\mathcal{B}_{1}\otimes\mathcal{B}_{2})$.
@@ 3.1 Normalizer matrices
Let $A=[a_{ij}]_{0\leq i\leq s-1,\,0\leq j\leq r-1}$ be an $s\times r$ matrix. For each $i\in\{0,1,\ldots,s-1\}$, we define the support of the $i$-th row of $A$ by
$Y_{i}:=\{\,j\in\{0,1,\ldots,r-1\}\mid a_{ij}\neq 0\,\}.$
We equip $Y_{i}$ with the natural order inherited from $\{0,1,\ldots,r-1\}$; that is, if
$Y_{i}=\{x_{1},x_{2},\ldots,x_{n}\},$
then
$x_{1}<x_{2}<\cdots<x_{n}.$
@@
Conversely, assume that the family $\{Y_{i}\}$ forms a partition of $\{0,1,\dots,s-1\}$. Let $a_{ij}=0$ and $a_{kj}\neq 0$ for some column $j$. Then $j\in Y_{k}\setminus Y_{i}$. If there exists a column $l$ such that $a_{il}\neq 0$ and $a_{kl}\neq 0$, then $l\in Y_{i}\cap Y_{k}$, which contradicts the assumption that the row supports form a partition. Hence, $a_{kl}=0$ as required. $\Box$
@@
An irredundant matrix $A=(a_{ij})$ is called a normalizer matrix if it satisfies the following conditions:
@@
    The row supports form a partition of the index set $\{0,1,\ldots,s-1\}$, where two indices $x$ and $y$ are equivalent if and only if
@@
Proof: Let
$i\colon\mathcal{B}\longrightarrow\mathcal{A}$
denote the inclusion map corresponding to the inclusion matrix $A$. Explicitly, this inclusion is given by
@@
Similarly, consider the inclusion
$\tilde{i}\colon\widetilde{\mathcal{B}}\longrightarrow\widetilde{\mathcal{A}}$
corresponding to the inclusion matrix $PAQ$, given by
@@
Define a *-isomorphism
$\theta\colon\mathcal{A}\longrightarrow\widetilde{\mathcal{A}}$
by
@@
$\theta\!\left(\bigoplus_{i=0}^{s-1}\mathrm{bl\text{-}diag}\bigl(X_{\tau(0)}\otimes\mathbf{1}_{a_{i\tau(0)}},\ldots,X_{\tau(r-1)}\otimes\mathbf{1}_{a_{i\tau(r-1)}}\bigr)\right)$ $=\bigoplus_{i=0}^{s-1}\mathrm{bl\text{-}diag}\bigl(X_{\tau(0)}\otimes\mathbf{1}_{a_{\sigma(i)\tau(0)}},$
$\hskip 79.6678pt\ldots,X_{\tau(r-1)}\otimes\mathbf{1}_{a_{\sigma(i)\tau(r-1)}}\bigr).$
@@
where each diagonal block $A_{ii}$ has all entries in every row equal and nonzero, and all off-diagonal blocks are zero.
Proof: Let $A$ be an normalizer matrix, and let $\{R_{k}\}_{k=1}^{p}$ denote the partition of $\{0,1,\ldots,s-1\}$ induced by the equivalence relation on row indices described above.
@@ 3.2 Normalizers of Subalgebras of Matrix Algebras
Throughout, let
$B=\bigoplus_{j=0}^{r-1}M_{m_{j}}(\mathbb{C})\subset M_{n}(\mathbb{C})=A$
be an inclusion of finite-dimensional von Neumann algebras, with inclusion matrix
$A=\begin{bmatrix}a_{0}&a_{1}&\cdots&a_{r-1}\end{bmatrix}$,
so that
$n=\sum_{j=0}^{r-1}m_{j}a_{j}.$
@@
Proof: Let the inclusion matrix for inclusion $\mathcal{B}=\mathbb{C}^{r}\subset M_{n}(\mathbb{C})$ be $\begin{bmatrix}a_{0}&a_{1}&\cdots&a_{r-1}\end{bmatrix}$ and let
$Q_{0},Q_{1},\ldots,Q_{r-1}$
be the minimal central projections of $\mathcal{B}$, where
@@
If inclusion $\mathcal{B}\subseteq\mathcal{A}$ is regular, then for every central projection $P\in Z(\mathcal{A})$, the induced inclusion
$P\mathcal{B}\subseteq P\mathcal{A}$
is also regular.
@@
For the second row, the nonzero entries occur at $j=2,3$, each equal to $2$. Thus
@@
A finite-dimensional $C^{*}$-algebra $\mathcal{A}$ is called homogeneous if there exist integers $r,n\geq 1$ such that
@@
Equivalently, $\mathcal{A}$ is a finite direct sum of simple matrix algebras, all of the same dimension.
@@
Let $\mathcal{B}\subset\mathcal{A}$ be a finite-dimensional inclusion as in (2.4), with inclusion matrix
$A_{\mathcal{B}}^{\mathcal{A}}=[a_{ij}]_{0\leq i\leq s-1,\;0\leq j\leq r-1}.$
Then the inclusion matrix of the induced inclusion
$\mathcal{Z}(\mathcal{B})\subset\mathcal{A}$
is given by
$A_{\mathcal{Z}(\mathcal{B})}^{\mathcal{A}}=\bigl[\,m_{j}\,a_{ij}\,\bigr]_{0\leq i\leq s-1,\;0\leq j\leq r-1}.$
@@
[…]
$\bigoplus_{j=0}^{r-1}\alpha_{j}\,\mathbf{1}_{m_{j}}\ \longmapsto\ \bigoplus_{i=0}^{s-1}\;\text{bl-diag}(\alpha_{0}\,\mathbf{1}_{m_{0}a_{i0}},\alpha_{1}\,\mathbf{1}_{m_{1}a_{i1}},\cdots,\alpha_{r-1}\,\mathbf{1}_{m_{r-1}a_{i,r-1}}).$
@@
Step 2: Regularity of the center forces equality of scalar multiplicities. Let $\mathcal{Z}(\mathcal{B})\cong\bigoplus_{j=0}^{r-1}\mathbb{C}\mathbf{1}_{m_{j}}$, and let its inclusion matrix in $M_{n}(\mathbb{C})$ be
$m_{0}a_{0}\;\;m_{1}a_{1}\;\;\cdots\;\;m_{r-1}a_{r-1}].$
By Proposition 3.7, regularity implies
@@
Since the products $m_{i}a_{i}$ are equal, it follows that all $a_{i}$ are also equal. Consequently, $\mathcal{B}$ is homogeneous, and all entries of the inclusion matrix are equal.
This completes the proof. $\Box$

### v2

@@ 1 Introduction
Normalizers occupy a central position in the structure theory of operator algebras, reflecting internal symmetries of inclusions and governing their associated crossed-product and groupoid constructions. Their importance is already evident in Dixmier’s seminal classification of maximal abelian subalgebras (MASAs) into singular, semi-regular, and regular (Cartan) types, where regularity is defined by generation of the ambient algebra by the normalizer. This perspective was further developed by Feldman and Moore [10], who showed that Cartan subalgebras give rise to measured equivalence relations, and later, Renault ([16]) discovered a $C^{*}$-algebraic analogue of this result in the language of groupoids.
In subfactor theory, normalizers play an important role (see, for instance, [12, 6, 2, 3, 15, 7]). Regular subfactors form a distinguished class encompassing crossed-product inclusions by finite groups and are closely related to depth-two phenomena and quantum symmetry. In particular, finite-index regular subfactors are known to admit unitary Pimsner–Popa bases and to arise from actions of weak Hopf algebras or quantum groupoids (see [3, 9]). Related classification results have also been obtained in [4] for finite-index regular inclusions of simple unital $C^{*}$-algebras, where regularity forces depth two and a crossed-product-type structure.
@@
Our approach is matrix-theoretic in nature. Recall that any inclusion $\mathcal{B}\subset\mathcal{A}$ of finite-dimensional von Neumann algebras is completely determined, up to isomorphism, by its inclusion matrix together with the dimension vectors of the simple summands. We introduce a new combinatorial invariant of an inclusion, called the normalizer matrix (see 3.2), which captures the structure of the normalizer purely at the level of the inclusion matrix. This invariant isolates the precise combinatorial obstructions to regularity and allows us to characterize regular inclusions in terms of elementary matrix data.
@@
As a consequence, any regular inclusion is isomorphic to tensor product and direct sum of inclusions $M_{k}(\mathbb{C})\subset M_{k}(\mathbb{C}),$ $\mathbb{C}\subset\oplus_{i=0}^{v-1}M_{l_{i}}(\mathbb{C})$ and $\mathbb{C}^{t}\subset M_{t}(\mathbb{C})$.
A second motivation for this work comes from the theory of unitary orthonormal (Pimsner–Popa) bases. Such bases play a central role in subfactor theory [14] and quantum information theory, where they are closely related to unitary error bases and to protocols such as teleportation and dense coding. While the existence of unitary orthonormal bases is automatic for several basic inclusions, a complete characterization for general finite-dimensional inclusions has remained open. In a seminal work, Werner [18] established a one-to-one correspondence between tight teleportation schemes for tripartite systems $M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ and orthonormal bases of unitaries in $M_{n}(\mathbb{C})$. This correspondence was subsequently extended by Conlon et al. [8] to inclusions of the form $N\subset M_{n}(\mathbb{C})$, demonstrating how unitary orthonormal bases on general multi-matrix algebras can be used to study quantum teleportation schemes and quantum graphs. In these settings, tight teleportation schemes are shown to arise from orthonormal unitary Pimsner–Popa bases, generalizing Werner’s original framework. In all such applications, it is essential that the unitary orthonormal basis be contained in the normalizer.
@@
Theorem B: (See Theorem 4.2, Corollary 4.3 and 4.4) Let $(\mathcal{B}\subset\mathcal{A},E)$ be an inclusion of finite-dimensional von Neumann algebras, where $E$ denotes the unique conditional expectation preserving the trace given in equation 4.8. Then the following statements are equivalent:
@@
Finally, we show that every regular inclusion has depth two, linking our results to (weak) Hopf-algebraic structures. More precisely, if the inclusion $\mathcal{B}\subset\mathcal{A}$ is regular, then the corresponding inclusion matrix has depth $2$ in the sense of [5].To see the connection between depth $2$ in subfactor theory and regularity, we refer to [13],[3]. Thus, our results may be viewed as a finite-dimensional analogue of the structure theory of regular inclusions developed in the subfactor and $C^{*}$-algebraic settings.
@@ 2 Preliminaries
Notation: The block diagonal form $\begin{bmatrix}A_{0}&\\ &\ddots\\ &&A_{n}\end{bmatrix}$ will be denoted by $\text{bl-diag}(A_{1},\cdots,A_{n})$.
@@
[…]
@@
[1] Let $(\mathcal{B}\subseteq\mathcal{A},E)$ be a subalgebra system, with $Am^{\prime}=n^{\prime}$, where $A$ is the inclusion matrix, and $m^{\prime},n^{\prime}$ are dimension vectors. Suppose $(\mathcal{B}\subseteq\mathcal{A},E)$ admits a unitary orthonormal basis with $d$-unitary elements. Then $A^{t}n^{\prime}=dm^{\prime}$.
@@
where $\mathcal{U(A)}$ denotes the group of unitaries of $\mathcal{A}$.Equivalently, the normalizer can also be defined as
@@
where $Ad_{U}(x)=UxU^{*}.$ The inclusion $\mathcal{B}\subseteq\mathcal{A}$ is called regular if $\mathcal{A}=\text{span}\;\mathcal{N}_{\mathcal{A}}(\mathcal{B})$.
@@
    Scalar Inclusion: $(\mathbb{C}\subset\oplus_{i=0}^{s-1}M_{n_{i}}(\mathbb{C}),E_{1})$ where $E_{1}$ is the unique conditional expectation preserving the state
    $\varphi(\oplus X_{i})=\frac{1}{\sum_{i=0}^{s-1}n_{i}^{2}}\sum_{i=0}^{s-1}n_{i}~\mbox{trace}~(X_{i}),~~X_{i}\in M_{n_{i}},0\leq i\leq(s-1).$ (2.5)
@@
Suppose $(\mathcal{B}_{i}\subseteq\mathcal{A}_{i},E_{i})$ for $i=1,2$ are two finite dimensional inclusions with $U$-property. Suppose $\{U_{1}(i),\ldots,U_{d_{i}}(i)\}$ is a unitary o.n.b. in $\mathcal{N}_{\mathcal{A}_{i}}(\mathcal{B}_{i})$ for $(\mathcal{B}_{i},\mathcal{A}_{i},E_{i})$ for $i=1,2.$ Take $\mathcal{B}=\mathcal{B}_{1}\otimes\mathcal{B}_{2}$. Consider it as a subalgebra of $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ with the inclusion matrix being the tensor product of inclusion matrices. Then it is seen easily that $E=E_{1}\otimes E_{2}$ is a conditional expectation map from $\mathcal{A}$ to $\mathcal{B}.$ For this inclusion, $(\mathcal{B}_{1}\otimes\mathcal{B}_{2}\subseteq\mathcal{A}_{1}\otimes\mathcal{A}_{2},E_{1}\otimes E_{2})$, we can observe that $\{U_{j}(1)\otimes U_{k}(2):0\leq j\leq d_{1};0\leq k\leq d_{2}\}$ is a unitary orthonormal basis in $\mathcal{N}_{\mathcal{A}_{1}\otimes\mathcal{A}_{2}}(\mathcal{B}_{1}\otimes\mathcal{B}_{2})$ as $\mathcal{N}_{\mathcal{A}_{1}}(\mathcal{B}_{1})\otimes\mathcal{N}_{\mathcal{A}_{2}}(\mathcal{B}_{2})\subseteq\mathcal{N}_{\mathcal{A}_{1}\otimes\mathcal{A}_{2}}(\mathcal{B}_{1}\otimes\mathcal{B}_{2})$.
@@ 3.1 Normalizer matrices
Let $A=[a_{ij}]_{0\leq i\leq(s-1),\,0\leq j\leq(r-1)}$ be an $s\times r$ matrix. For each $i\in\{0,1,\ldots,(s-1)\}$, we define the support of the $i$-th row of $A$ by
$Y_{i}:=\{\,j\in\{0,1,\ldots,(r-1)\}\mid a_{ij}\neq 0\,\}.$
We equip $Y_{i}$ with the natural order inherited from $\{0,1,\ldots,r-1\}$; that is, if $Y_{i}=\{x_{1},x_{2},\ldots,x_{n}\},$ then $x_{1}<x_{2}<\cdots<x_{n}.$
@@
Conversely, assume that the family $\{Y_{i}\}$ forms a partition of $\{0,1,\dots,(s-1)\}$. Let $a_{ij}=0$ and $a_{kj}\neq 0$ for some column $j$. Then $j\in Y_{k}\setminus Y_{i}$. If there exists a column $l$ such that $a_{il}\neq 0$ and $a_{kl}\neq 0$, then $l\in Y_{i}\cap Y_{k}$, which contradicts the assumption that the row supports form a partition. Hence, $a_{kl}=0$ as required. $\Box$
@@
An irredundant matrix $A=[a_{ij}]$ is called a normalizer matrix if it satisfies the following conditions:
@@
    The row supports form a partition of the index set $\{0,1,\ldots,(s-1)\}$, where two indices $x$ and $y$ are equivalent if and only if
@@
Proof: Let $i\colon\mathcal{B}\longrightarrow\mathcal{A}$ denote the inclusion map corresponding to the inclusion matrix $A$. Explicitly, this inclusion is given by
@@
Similarly, consider the inclusion $\tilde{i}\colon\widetilde{\mathcal{B}}\longrightarrow\widetilde{\mathcal{A}}$ corresponding to the inclusion matrix $PAQ$, given by
@@
Define a *-isomorphism $\theta\colon\mathcal{A}\longrightarrow\widetilde{\mathcal{A}}$ by
@@
$\theta\!\left(\bigoplus_{i=0}^{s-1}\mathrm{bl\text{-}diag}\bigl(X_{\tau(0)}\otimes\mathbf{1}_{a_{i\tau(0)}},\ldots,X_{\tau(r-1)}\otimes\mathbf{1}_{a_{i\tau(r-1)}}\bigr)\right)=\bigoplus_{i=0}^{s-1}\mathrm{bl\text{-}diag}$ $\bigl(X_{\tau(0)}\otimes\mathbf{1}_{a_{\sigma(i)\tau(0)}},\ldots,$
$X_{\tau(r-1)}\otimes\mathbf{1}_{a_{\sigma(i)\tau(r-1)}}\bigr).$
@@
where each diagonal block $A_{ii}$ has the property that all entries in every row are equal and nonzero, and all off-diagonal blocks are zero.
Proof: Let $A$ be a normalizer matrix, and let $\{R_{k}\}_{k=1}^{p}$ denote the partition of $\{0,1,\ldots,s-1\}$ induced by the equivalence relation on row indices described above.
@@ 3.2 Normalizers of Subalgebras of Matrix Algebras
Proof: Let the inclusion matrix for inclusion $\mathcal{B}=\mathbb{C}^{r}\subset M_{n}(\mathbb{C})$ be $\begin{bmatrix}a_{0}&a_{1}&\cdots&a_{r-1}\end{bmatrix}$ and let $Q_{0},Q_{1},\ldots,Q_{r-1}$ be the minimal central projections of $\mathcal{B}$, where
@@
If inclusion $\mathcal{B}\subseteq\mathcal{A}$ is regular, then for every central projection $P\in\mathcal{Z(A)}$, the induced inclusion $P\mathcal{B}\subseteq P\mathcal{A}$ is also regular.
@@
For the second row, the nonzero entries occur at $j=2,3$ each equal to $2$. Thus
@@
A finite-dimensional von Neumann algebra $\mathcal{A}$ is called homogeneous if there exist integers $r,n\geq 1$ such that
@@
Equivalently, $\mathcal{A}$ is a finite direct sum of matrix algebras, all of the same dimension.
@@
Let $\mathcal{B}\subset\mathcal{A}$ be a finite-dimensional inclusion as in (2.4), with inclusion matrix $A_{\mathcal{B}}^{\mathcal{A}}=[a_{ij}]_{0\leq i\leq(s-1),\;0\leq j\leq(r-1)}.$ Then the inclusion matrix of the restricted inclusion $\mathcal{Z}(\mathcal{B})\subset\mathcal{A}$ is given by
$A_{\mathcal{Z}(\mathcal{B})}^{\mathcal{A}}=\bigl[\,m_{j}\,a_{ij}\,\bigr]_{0\leq i\leq(s-1),\;0\leq j\leq(r-1)}.$
@@
$i:\ \bigoplus_{j=0}^{r-1}\mathbb{C}\,\mathbf{1}_{m_{j}}\$$\longrightarrow\ \bigoplus_{i=0}^{s-1}M_{n_{i}}$
$\bigoplus_{j=0}^{r-1}\alpha_{j}\,\mathbf{1}_{m_{j}}\$$\longmapsto\ \bigoplus_{i=0}^{s-1}\;\text{bl-diag}(\alpha_{0}\,\mathbf{1}_{m_{0}a_{i0}},\alpha_{1}\,\mathbf{1}_{m_{1}a_{i1}},\cdots,\alpha_{r-1}\,\mathbf{1}_{m_{r-1}a_{i(r-1)}}).$
@@
Step 2: Regularity of the center forces equality of scalar multiplicities. Let $\mathcal{Z}(\mathcal{B})\cong\bigoplus_{j=0}^{r-1}\mathbb{C}\mathbf{1}_{m_{j}}$, and its inclusion matrix in $M_{n}(\mathbb{C})$ is $m_{0}a_{0}\;\;m_{1}a_{1}\;\;\cdots\;\;m_{r-1}a_{r-1}].$ By Proposition 3.7, regularity implies
@@
Since the products $m_{i}a_{i}$ are equal, it follows that all $a_{i}$ are also equal. Consequently, $\mathcal{B}$ is homogeneous, and all entries of the inclusion matrix are equal. $\Box$